### v1

# Computational Modelling of Thixotropic Multiphase Fluids

Abstract
Multiphase systems are ubiquitous in engineering, biology, and materials science, where understanding their complex interactions and rheological behavior is crucial for advancing applications ranging from emulsion stability to cellular phase separation. This study presents a numerical methodology for modeling thixotropic multiphase fluids, emphasizing the transient behavior of viscosity and the intricate interactions between phases. The model incorporates phase-dependent viscosities, interfacial tension effects, and the dynamics of phase separation, coalescence, and break-up, making it suitable for simulating systems with complex flow regimes. A key feature of the methodology is its ability to capture thixotropic behavior, where viscosity evolves over time due to microstructural changes induced by shear history. This approach enables the simulation of aging and recovery processes in materials such as gels, emulsions, and biological tissues. The model is rigorously validated against benchmark cases, demonstrating its accuracy in predicting multiphase systems under static and dynamic conditions. Subsequently, the methodology is applied to investigate systems with varying levels of microstructural evolution, revealing the impact of thixotropic dynamics on overall system behavior. The results provide new insights into the time-dependent rheology of multiphase fluids and highlight the versatility of the model for applications in industrial and biological systems involving complex fluid interactions.
keywords:
Smoothed Dissipative Particle Dynamics (SDPD) , Multiphase flows , Liquid-Liquid Phase Separation , Thixotropy
journal: Computer Physics Communications
[…]
[label1]organization=Escuela de Ingeniería de Bilbao, Universidad del País Vasco (EHU),addressline=Plaza Ingeniero Torres Quevedo, 1, city=Bilbao, postcode=48013, state=Basque Country, country=Spain
[…]
[label2]organization=Basque Center for Applied Mathematics (BCAM),addressline=Alameda de Mazarredo, 14, city=Bilbao, postcode=48400, state=Basque Country, country=Spain
[…]
[label3]organization=Basque Foundation for Science (IKERBASQUE),addressline=Calle de Maria Diaz de Haro 3, city=Bilbao, postcode=480013, state=Basque Country, country=Spain
[…]
[label4]organization=Complex Fluids Research Group, Department of Chemical Engineering, Faculty of Science and Engineering, Swansea University,addressline= SA1 8EN, city=Swansea, country=UK
1 Introduction
Multiphase systems are crucial to engineering, biology, and materials science, where the existence of complex microstructural dynamics often lead to non-trivial rheological behavior, such as shear-thinning or thixotropic properties. In industrial applications, understanding emulsion and foam stability is critical for improving product quality and efficiency [63]. In biology, liquid-liquid phase separation within cells organizes cellular components and influences diseases like neurodegenerative disorders [19, 10, 21, 14]. Similarly, interactions among blood components, such as plasma and red blood cells, are essential for processes like clotting under varying physiological conditions [4, 9, 2]. In materials science, understanding phase interactions is crucial for designing smart materials, composites, and alloys with desired properties [64, 65]. Accurately modeling these systems, including phase separation dynamics and time-dependent material properties [62], remains challenging, driving the development of advanced grid-based and mesh-free methods.
Grid-based methods such as the Finite Element [77, 78, 79], Lattice Boltzmann Method (LBM) [80, 81, 82], methods based on the Volume of Fluid (VOF) [72, 73, 71], Level Set (LS) [74, 75, 76] or Immersed Boundary Method (IBM) [97, 98], are widely used for tracking interfaces and capturing phase interactions. These methods are particularly effective for problems involving sharp interfaces. However, they can struggle with dynamic interface tracking, requiring significant computational resources for high-resolution simulations. Additionally, spurious non-physical velocity fields near the interfaces (resulting from discretization errors) can change the shape or stability of the simulated phases [66, 67, 68]. Mesh-free methods, such as Smoothed Particle Hydrodynamics (SPH) [25, 94, 95, 6] and Dissipative Particle Dynamics (DPD) [83, 96, 84, 85], have gained popularity for their ability to handle complex geometries and dynamic interfaces without the need for explicitly tracking the interfaces. These methods are particularly advantageous for simulating multiphase flows with large deformations[6] and complex interactions[87]. However, they can suffer from numerical instabilities, especially at interfaces, where large differences in viscosity and density can generate pressure oscillations and unphysical artifacts, such as gaps or particle mixing at the interface [8, 3, 7, 70, 5]. Additionally, implementing accurate boundary conditions at walls and interfaces for fluids with varying viscosities is complex and can result in errors affecting flow behavior near these regions.
Smoothed dissipative particle dynamics (SDPD) [12, 11] is a mesoscopic technique that combines the discretization scheme of SPH and the consistent thermal fluctuations of DPD, effectively reproducing the fluctuating Navier-Stokes equations. SDPD has been successfully applied to model a range of synthetic [60, 20, 37] and biological systems [18, 26, 27]. Recently, Lei et al. [17] developed a SDPD scheme suitable to model multiphase systems incorporating surface tension effects via a pairwise interaction forces. This scheme was able to capture complex phenomena such as droplet coalescence and interface stabilization in the presence of thermal noise, showing results consistent with theoretical predictions and experimental data. The method’s ability to incorporate thermal fluctuations makes it particularly suitable for studying micro to nanoscale systems where thermal fluctuations can play a role in phase interactions and dynamics. However, the application of SDPD to multiphase flows with complex rheological behavior, such as thixotropic materials, which is crucial for accurately simulating complex fluids like gels or emulsions that exhibit microstructural changes during flow, remains an emerging field of research.
In this work, we introduce a comprehensive SDPD model designed to address the challenges inherent in modeling multiphase systems with complex rheological behavior. The proposed model incorporates several key features to accurately capture the dynamics of such systems. First, it incorporates interfacial tension effects, which are essential for understanding interactions at phase boundaries, such as those observed in emulsions. Second, the model accounts for the dynamics of phase separation, including phenomena like coalescence and break-up, which are critical for predicting the stability and structural evolution of multiphase systems under varying flow conditions. Third, it explicitly accounts for phase-dependent viscosities, enabling precise representation of systems where fluids or materials exhibit distinct flow properties. Additionally, the methodology integrates a thixotropic viscosity model to capture time-dependent rheological behavior, allowing for the simulation of aging and recovery processes in materials such as gels and emulsions. By incorporating microstructural dynamics, the model provides a versatile framework for studying systems that undergo structural changes during flow.
The present manuscript is organized as follows. In the first section, we introduce our methodology and numerical model. In the second section, we validate the robustness and accuracy of our implementation for both Newtonian and Non-Newtonian (thixotropic) fluids in static and dynamic states. We further illustrate the flexibility of the methodology for various applications in biology and microfluidics, including liquid-liquid phase separation, thixotropic emulsions flow, and complex-microfluidic geometries to control merging and splitting in multiphase flows. Finally, we provide the main conclusions and recommendations arising from this work.
2 Numerical methodology
In SDPD, a fluid is discretized using particles with a volume $V_{i}$, leading to a particle number density $d_{i}=1/V_{i}=\sum_{j}W(r_{ij},h)$. Where $W(r_{ij},h)$ is a kernel function that depends on the distance $r_{ij}=|{\bf{r}}_{i}-{\bf{r}}_{j}|$ between particles $i$ and $j$, has a finite support $h$ and is normalized to one. The evolution of the particle position is given by the equation ${\text{d}{\bf{r}}_{i}}/{\text{d}t}={\bf v}_{i}$, where ${\bf v}_{i}$ is the velocity of the $i$-th particle. The momentum’s stochastic differential equation is given by
$m_{i}\frac{\text{d}{\bf v}_{i}}{\text{d}t}$ $=-\sum_{j}\left[\frac{p_{i}}{d_{i}^{2}}+\frac{p_{j}}{d_{j}^{2}}\right]F_{ij}{\bf r}_{ij}$ (1)
$-\sum_{j}\left[a_{ij}{\bf v}_{ij}+b_{ij}({\bf v}_{ij}\cdot{{\bf e}_{ij}}){{\bf e}_{ij}}\right]\frac{F_{ij}}{d_{i}d_{j}}$
$+\sum_{j}\left(A_{ij}\text{d}{\bf\overline{W}}_{ij}+B_{ij}\frac{1}{D}\text{tr}[\text{d}{\bf{W}}_{ij}]\right)\cdot\frac{{{\bf e}_{ij}}}{\text{d}t},$ (2)
where $p_{i}$ and $p_{j}$ are the pressure of particles $i$ and $j$, respectively. The pressure of each particle is determined by a suitable density-dependent equation of state $p=f(\rho)$, as discussed later. The term $m_{i}=\rho\Delta^{D}$, denotes the mass of the particle that depends on the density $\rho$, the equilibrium interparticle distance $\Delta$ and the dimension of the system $D$. The term $F_{ij}=-\nabla W(r_{ij},h)/r_{ij}$ is a positive function introduced to account for the interaction forces between particles. Here, we adopt a Lucy’s Quartic kernel [28] of the form $W(r)=\mathrm{w_{0}}(1+3\frac{r}{h})(1-\frac{r}{h})^{3}$ if $r/h\leq 1$ and $W(r)=0$ if $r/h>1$, where $\mathrm{w_{0}}=5/(\pi h^{2})$ for two dimensions and $\mathrm{w_{0}}=105/(16\pi h^{3})$ for three dimensions. The second term in (2) corresponds to viscous contributions, where ${\bf v}_{ij}={\bf v}_{i}-{\bf v}_{j}$ is the relative velocity between particles, ${{\bf e}_{ij}}={\bf r}_{ij}/|{\bf r}_{ij}|$ is the unit vector, $a_{ij}$ and $b_{ij}$ are friction coefficients related to the pairs shear viscosity ($\eta_{ij}$) and bulk viscosity ($\zeta_{ij}=\eta_{ij}(2D-1)/D$), with $a_{ij}={(D+2)\eta_{ij}}/{D}-\zeta_{ij}$ and $b_{ij}=(D+2)(\zeta_{ij}+{\eta_{ij}}/{D})$. Here, we consider the viscosity between particles $i$ and $j$ expressed as $\eta_{ij}=2(\eta_{i}\eta_{j})/(\eta_{i}+\eta_{j})$ The last term in (2) incorporates thermal fluctuations into the momentum balance by including a matrix of independent increments of the Wiener process, $d{\bf{W}}_{ij}$ and its traceless symmetric part $d{\bf\overline{W}}_{ij}$ given by
$d{\bf\overline{W}}_{ij}=\frac{1}{2}[d{\bf{W}}_{ij}+d{\bf{W}}_{ij}^{T}]-\frac{\delta^{\alpha\beta}}{D}\text{tr}[d{\bf{W}}_{ij}].$ (3)
The parameters $A_{ij}$ and $B_{ij}$ are the amplitude of the noise, related to the friction coefficients ($a_{ij}$ and $b_{ij}$) and is given by
$A_{ij}$ $=\left[4\ k_{B}T\ a_{ij}\ \frac{F_{ij}}{d_{i}d_{j}}\right]^{-1/2}$ (4)
$B_{ij}$ $=\left[4\ k_{B}T\ \left(b_{ij}-a_{ij}\frac{D-2}{D}\right)\ \frac{F_{ij}}{d_{i}d_{j}}\right]^{-1/2},$ (5)
where $k_{b}$ is the Boltzmann constant and $T$ temperature.
2.1 Thixotropic model
When exposed to continuous shear stress, some non-Newtonian fluids can exhibit a time-dependent shear-thinning or thixotropic behavior, which is characterized by a gradual decrease in viscosity followed by a recovery upon removal of the stress. To capture this behavior, we can follow the phenomenological description given by Le-Cao et al. [16], where viscosity of a fluid is given by $\eta(t,x)=\eta_{\infty}\left[1+\alpha f(t,\dot{\gamma}(x))\right]$, where $\eta(t)$ is the time-dependent shear viscosity, $\eta_{\infty}$ is the limiting viscosity, $\alpha$ is a constitutive constant that determines the maximum viscosity and $f=[0:1]$, as indicated by Rossi et al. [62], is a microstructural scalar parameter describing the current state of the microstructure from fully destroyed ($f=0$) to completely developed ($f=1$). The time evolution of $f$ is given by $\dot{f}=a-(a+b\dot{\gamma})f$, where $a$ is a constant that represents the rate of formation of the microstructure, and $b\dot{\gamma}$ the rate of destruction of the microstructure under shear conditions. The strain-rate tensor $\dot{\boldsymbol{\gamma}}$ is given by $\dot{\boldsymbol{\gamma}}=(\nabla{\bf v})+(\nabla{\bf v}^{T})$. The strain-rate tensor second invariant $\dot{\gamma}$ can be evaluated as:
$\dot{\gamma}=\sqrt{\frac{1}{2}\it\Pi_{\dot{\gamma}}}=\left[\frac{1}{2}\{\dot{\boldsymbol{\gamma}}:\dot{\boldsymbol{\gamma}}\}\right]^{1/2}$ (6)
The velocity gradient tensor $\nabla{\bf v}$ is given by
$(\nabla{\bf v})^{\alpha\beta}_{i}=\sum_{j}\frac{({\bf v}^{\alpha}_{i}-{\bf v}^{\alpha}_{j})({\bf r}_{i}^{\beta}-{\bf r}_{j}^{\beta})}{d_{j}r_{ij}}\frac{\partial W(r_{ij},h)}{\partial r},$ (7)
where $\alpha,\beta$ are the column and row indices in the matrix tensor and $\partial W(r_{ij},h)/\partial r$ the gradient of the kernel. By considering the ratio between the formation and destruction and the characteristic time for microstructure formation, we can introduce the parameter $\beta=b/a$ and the thixotropic time scale $\lambda_{0}=1/a$. Accordingly, the temporal evolution can be rewritten as $\dot{f}=[1-(1+\beta\dot{\gamma})f]\lambda_{0}^{-1}$ and, after time integration, the scalar parameter $f$ can be calculated as
$f(t)=\frac{1}{1+\beta\dot{\gamma}}\left(1-\exp^{-\frac{(1+\beta\dot{\gamma})t}{\lambda_{0}}}\right)+f_{0}\exp^{-\frac{(1+\beta\dot{\gamma})t}{\lambda_{0}}},$ (8)
where $f_{0}$ is the initial value of the scalar parameter $f$. Similar to the viscosity between a pair of particles $i$ and $j$, any pair parameter $X_{ij}$ of the thixotropic model is approximated by $X_{ij}=2(X_{i}X_{j})/(X_{i}+X_{j})$. For instance, we have for the limiting viscosity $(\eta_{\infty})_{ij}=2((\eta_{\infty})_{i}(\eta_{\infty})_{j})/((\eta_{\infty})_{i}+(\eta_{\infty})_{j})$, and similarly for the rest of factors.
2.2 Multiphase Model
Interfacial tension can be included in the model using an additional pairwise interaction force $\mathbf{F}^{int}_{ij}$, to the momentum balance (2). This interaction force takes the form[25]
$\mathbf{F}^{int}_{ij}=\mathbf{F}^{int}(\mathbf{r}_{ij})=-s_{ij}\phi(r_{ij})(\mathbf{r}_{ij}/r_{ij}),$ (9)
where $s_{ij}$ and $\phi$ determine the strength and functional form of the potential, respectively. This force is short-range repulsive and long-range attractive [17]. For two different phases $\alpha$ and $\beta$ the factor $s_{ij}$ is related to the domains $\Omega_{\alpha}$ and $\Omega_{\beta}$ as
$s_{ij}=\begin{cases}s_{\alpha\alpha}&\text{if }\mathbf{r}_{i}\in\Omega_{\alpha}\text{ and }\mathbf{r}_{j}\in\Omega_{\alpha},\\ s_{\beta\beta}&\text{if }\mathbf{r}_{i}\in\Omega_{\beta}\text{ and }\mathbf{r}_{j}\in\Omega_{\beta},\\ s_{\alpha\beta}&\text{if }\mathbf{r}_{i}\in\Omega_{\alpha}\text{ and }\mathbf{r}_{j}\in\Omega_{\beta},\end{cases}$ (10)
where
$s_{\alpha\alpha}=s_{\beta\beta}=10^{k}s_{\alpha\beta}=\frac{1}{1-10^{-k}}n^{-2}_{eq}\frac{\sigma_{0}}{\lambda_{m}}$ (11)
Here, $k$ is a constant that must be greater than 1, $\sigma_{0}$ is a macroscopic surface tension. $n_{eq}=\Delta^{-D}$ is the equilibrium particle density which depends on the grid size $\Delta$. The factors $\lambda_{m}$ and $\phi(r_{ij})$ are given by
$\lambda_{m}$ $=\sqrt{2\pi}(-A\epsilon_{0}^{5}+\epsilon^{5}),$ (12)
$\phi(r_{ij})$ $=r_{ij}\left[-Ae^{-\frac{r^{2}_{ij}}{2\epsilon_{0}^{2}}}+e^{-\frac{r^{2}_{ij}}{2\epsilon^{2}}}\right],$ (13)
where $A$ represents the magnitude of the ratio between repulsive and attractive forces and $\epsilon$ and $\epsilon_{0}$ are compact-support dependent functions. Following Tartakovsky and Panchenko [25], we use $\epsilon=h/3.5$, $\epsilon_{0}=\epsilon/2$ and $A=(\epsilon/\epsilon_{0})^{3}$. This choice ensures that the virial pressure ($P_{v}=-\pi(-A\epsilon_{0}^{4}+\epsilon^{4})n_{eq}^{2}s_{\alpha\alpha}$) is negative, which is necessary for the stability of the interfaces [25]. However, we must note, that the presence of a negative virial pressure can also lead to numerical instabilities in the system, especially when the system is not in equilibrium. In order to prevent this viral pressure affects the global system pressure, we introduce an equation of state fo the form
$P=\frac{c_{s}^{2}\rho_{0}}{7}\left[\left(\frac{\rho}{\rho_{0}}\right)^{7}-1\right]+P_{b},$ (14)
where $\rho_{0}$ is the equilibrium density for the system, $c_{s}$ is the speed of sound, and $P_{b}=-P_{v}=-\pi(-A\epsilon_{0}^{4}+\epsilon^{4})n_{eq}^{2}s_{\alpha\alpha}$ is a background pressure that ensures $P>0$. This component balance the system and avoid spurious effects of global negative pressure. It is important to note that the presence of this background pressure can also influence the dynamics of the system, especially in flow conditions. In the Results section, we show that the choice of $P_{b}$ does not introduce artifacts into the simulations, under different dynamic conditions.
In our simulations, we adopt the well-known Velvet-velocity algorithm [36] for the temporal evolution of the position and velocity of the particles. We use a time step $\Delta t$ that ensure numerical stability[37], satisfying both: $\Delta t\leq 0.125h^{2}\rho/\eta$ (the viscous diffusion criteria [38]) and $\Delta t\leq 0.25h/c_{s}$ (the Courant-Friedrichs-Lewy (CFL) condition [39]).
3 Results and Discussion
3.1 Validation of the numerical methodology
We conduct a series of benchmark simulations to ensure the accuracy and robustness of our implementation. First addressing the correct description of the surface tension and contact angles in static and dynamic cases. Then, we validate the thixotropic model by simulating a transient viscosity case. The results of these simulations are compared with theoretical predictions and previous numerical results
3.1.1 Static validation: Surface tension and contact angles
(a) Pressure difference versus curvature $1/R$
(b) Retraction of a stretched droplet
(c) Static contact angle
Figure 1: Validation of the methodology for multiphase static cases of (a) droplet surface tension, (b) retraction of a stretched droplet and (c) static contact angle between different phases.
As several research studies have shown (see [22, 23, 24, 15, 25, 17]), within the validation of numerical methods for multiphase flows, the analysis of the surface tension play an important role. The Young Laplace equation indicates that the difference in pressure between the droplet ($P_{d}$) and the external fluid ($P_{f}$) is directly proportional to the surface tension ($\sigma$) and inversely proportional to the radius ($R$), i.e., $P_{d}-P_{f}=2\sigma/R$. We model the effect of surface tension starting with a square-shaped particle array, which evolves in time to a circular droplet. The size of the initial square varies to produce a range of equilibrium radius. For these simulations, we use $\sigma=0.25$ and $\sigma=1$, with an equilibrium particle density $n_{eq}=25$ and $k_{b}T\to 0$. In Figure 1(a), we present the evolution of the pressure difference as a function of curvature ($1/R$). Consistent with the Young-Laplace equation, the slopes obtained from the linear fitting, $0.53$ and $2.08$ corresponds to $\sim 2\sigma$. We further validate the consistency of our method, measuring the deformation of an initially elliptical droplet as it retracts to a circular shape due to interfacial tension effects[30, 31, 32, 33]. The initial ellipse is defined by the semi-major axis $A$ and the semi-minor axis $B$, with an initial Taylor deformation parameter[29] ${\mathcal{D}}_{t=0}=(A-B)/(A+B)$. The retraction process is governed by the balance between surface tension and viscous forces. The evolution of the of the Taylor parameter $\mathcal{D}$ can be described by[30, 31, 32]
$\mathcal{D}=\mathcal{D}_{t=0}\exp\left(-\frac{\sigma}{\eta R_{0}}\frac{40(\lambda+1)}{(2\lambda+3)(19\lambda+16)}t\right),$ (15)
where $R_{0}$ the final circle radius, $\lambda$ the ratio between the viscosity of the phases and $t$ the stretching time. Here, we use three different viscosity ratios $\lambda=[1,5,10]$, and use Principal Component Analysis (PCA) to compute the eigenvalues and eigenvectors of the droplets particles, providing the magnitude and orientation of the ellipse’s principal axes. In Figure 1(b), we compare our results for the time evolution of $\mathcal{D}$ with (15), evidencing a good agreement between the numerical and theoretical results.
We now validate the model’s ability to consistently capture contact angle variations. In (9), the strength factor $s_{xx}$ (with $xx$ denoting the phase pair, e.g., $A$, $B$, …, $n$), governs interfacial interactions and equilibrium configurations [25]. To illustrate, we consider a system of four phases where phase $1$ contains droplets of phases $2$, $3$, and $4$, initially non-contacting, as shown in detail in Figure 1(c). Interaction strengths are defined as $s_{11}=s_{22}=s_{33}=s_{44}$ and $s_{12}=s_{13}=s_{14}=s_{23}=s_{24}=s_{34}$. Under these conditions, the system evolves into an equilibrium configuration with a symmetric triple contact angle of $120^{\circ}$ between each pair of interfaces. Similarly, varying the interaction strengths should enable us to control the static contact angles on solid surfaces[15, 34]. In Figure 1(c), we present the contact angle variation between a fluid (1), a droplet (2) and a solid surface (3). To estimate the contact angle, we measure the droplet height $a_{d}$ and width $2b_{d}$, such that $\theta={\pi}/{2}-\arcsin({b_{d}^{2}-a_{d}^{2}}/{b_{d}^{2}+a_{d}^{2}})$ [25]. To obtain angles between $0^{o}$ and $90^{o}$ it is necessary that the interaction force between fluid-droplet be less than the interaction force between droplet-solid ($s_{13}<s_{23}$). In the other hand, for $90^{o}$ to $180^{o}$, it is necessary that the interaction force between fluid-droplet be greater than the interaction force between droplet-solid ($s_{13}>s_{23}$). We compare the measured angles between 0 and 90 plotted against the droplet-solid interaction force ($s_{23}$) normalized to its maximum value ($s_{23,max}$). In Figure 1(c), we show that the measured variation in the contact angle is consistenty with previously reported numerical results  [15, 34]. This static validation confirms that the implemented multiphase scheme reliably captures interfacial behavior and surface tension effects across fluid–fluid and fluid–solid boundaries.
3.2 Dynamic validation: Poiseuille and shear flow
As discussed in the description of the numerical methodology, the interaction force (Equation 9) induces negative virial pressure in the system, which can lead to numerical instabilities. To avoid this, we introduce a background pressure $P_{b}$ (directly linked to $\sigma_{0}$) in (14) that counterbalances the negative virial pressure. To verify that $P_{b}$ does not introduce artifacts, we initially evaluate single-phase flows under various configurations, then, we validate dynamic conditions for multiphase systems in Poiseuille and shear flows. For single-phase flows, the system’s behavior should align with the standard SDPD formulation (2), regardless of the added pairwise force or background pressure. A presents a comparison of velocity and stress profiles for single-phase fluid flows against theoretical predictions and prior numerical results. The analysis demonstrates that incorporating $P_{b}$ not only avoids introducing artifacts into the simulation but is also essential for maintaining system stability.
In the following, we validate the multiphase model by simulating two benchmark cases: a Poiseuille flow in a channel with two fluid phases and the dynamics of a droplet suspended in a liquid domain. The first case allows us to assess the model’s ability to capture the velocity profile and interface stability under varying viscosity ratios, while the second case tests the model’s capability to simulate droplet deformation and breakup under shear flow conditions. First, we consider a Poiseuille flow in a channel containing two fluid phases, $A$ and $B$, initially distributed in the lower and upper half of the channel, respectively. We set a constant viscosity for the phase $A$ ($\eta_{A}=1$) and different the viscosities for the phase $B$, by varying viscosity ratio $\lambda=\eta_{B}/\eta_{A}$ in the range $\{1,2,5,1\}$. For this setting, the theoretical velocity profile follows [24]
$v(x)=\rho gL^{2}\begin{cases}-\frac{1}{2\mu_{B}}\left(\frac{y}{L}\right)^{2}+a\frac{1}{\mu_{B}}\frac{y}{L}&0<y<L/2,\\ \frac{1}{2\mu_{A}}\left(1-\left(\frac{y}{L}\right)^{2}\right)+a\frac{1}{\mu_{A}}\left(\frac{y}{L}-1\right)&L/2\leq y<L,\end{cases}$ (16)
where
$a=\left(\frac{3}{4\mu_{A}}+\frac{1}{4\mu_{B}}\right)\left(\frac{1}{\mu_{B}}+\frac{1}{\mu_{A}}\right)^{-1},$ (17)
In Figure 2(a), we compare the velocity profiles simulated for the different viscosity ratios with the expected theoretical predictions. Notably, the model accurately capture the retraction in the parabolic profile as the viscosity ratio increases, preserving the stability of the interface between the two phases.
(a) Multiphase flow at different $\lambda$
(b) D vs Ca
(c) Droplet break-up
Figure 2: Validation of the methodology for multiphase Poiseuille flows using (a) a channel with two phases at different viscosity ratios, (b) Taylor deformation vs Capillary number for a droplet under a shear flow and (c) droplet break-up.
Next, we validate the model by simulating the dynamics of a droplet ($d$) of radius $r_{0}$, suspended in a continuous fluid ($f$) phase under shear flow conditions. The droplet is initially spherical and is subjected to a shear flow induced by a constant velocity gradient $\dot{\gamma}$. The capillary number $Ca=(\dot{\gamma}\ r_{0}\ \mu_{f})/\sigma$ quantify the relative effect of viscous forces to surface tension, where $r_{0}$ is the initial radius of the droplet, $\mu_{f}$ is the viscosity of the continuous phase.
In Figure 2(b), we present the measured Taylor deformation parameter $\mathcal{D}$ for different capillary numbers $Ca=\{0.05,0.1,0.2,0.3\}$. We consider a droplet of size $r_{0}=2.5$, within two walls moving at velocity $U$ and $-U$, respectively, in a square channel of size $L=6r_{0}$ The physical properties of the droplet and fluid are $\rho_{f}=\rho_{d}=1$, $\mu_{f}=\mu_{d}=5$ and $\sigma=2.5$. We vary $Ca$ by changing the walls velocity, leading to Reynolds number $Re=(\rho_{c}\ \dot{\gamma}\ r_{0}^{2})/\mu_{f}$ in the range $\{0.025,0.05,0.075,0.1\}$ always smaller than 1. In Figure 2(b), we can observe an excellent agreement of our numerical results with the theoretical predictions[41] for $Ca<0.4$, where $\mathcal{D}$ can be approximated as $D=(19\lambda+16)/(16\lambda+16)Ca$.
In Figure 2(c), we further explore the droplet dynamics as the capillary number increases, where the droplet undergoes breakup due to the unbalance between viscous forces and surface tension. Figure 2(c) shows the evolution from the undeformed state to the break-up of a droplet with $r_{0}=3$, $L=10r_{0}$, $\sigma=1$, $\mu_{f}=\mu_{d}=1$ and $\gamma=0.3$ corresponding to a $Ca=0.9$. The breakup occurs when the droplet’s aspect ratio exceeds a critical value, leading to the formation of smaller droplets. The breakup process is characterized by the formation of a neck at the droplet’s center, which eventually leads to the separation of the droplet into two smaller droplets. This behavior is consistent with previous studies on droplet dynamics in shear flows[41, 42, 43].
3.3 Thixotropic model validation
We validate our implementation of the thixotropic model considering a single-phase fluid in a simple shear flow within two parallel plates moving in opposite directions. For these simulations, we define a distance between plates $L=10$ with an initial interparticle distance $\Delta x=0.2$. The fluid properties are set as $f_{0}=1$, $\eta_{\infty}=15$, $\rho=1$, and $a=b=0.025$ (all within the adopted unit system), to fulfill the condition $\lambda_{0}>L^{2}\rho/\eta_{\infty}$ [62]. Following Rossi et al. [62], we track the evolution of the shear stress $\tau$ and the steady state viscosity $\eta$, from the measured tangential force on the walls $F_{x}$, such that $\tau={F_{x}}/{A}$ and $\eta(\dot{\gamma})={F_{x}}/{A\dot{\gamma}}$, where $A$ is the wall area.
In Figure 3, we summarize the results of the thixotropic model validation, evidencing the accuracy of the model in both steady and transient conditions. In Figure 3(a) and Figure 3(b), we present the variation in the measured steady state values of $\tau$ and $\eta$ (for various values of $\alpha=[1,2,4]$), as the shear rate $\dot{\gamma}$ increases, following an excellent agreement with the theoretical model Equation 8 (see Section 2). In Figure 3(c), we further present the transient behavior of $\eta(t)$ for a fluid with $\alpha=1$ under different shear rates. The results evidence the microstructural evolution of the fluid, where the viscosity decreases with time as the microstructure forms and grows at different rates. For the lowest values of shear rate, noticeable oscillations in the instantaneous force measurement lead to small fluctuation in the measured $\eta$. However, these numerical oscillations do not affect the overall trend of the viscosity evolution, which remains consistent with the expected behavior of a thixotropic fluid.
(a) Shear stress
(b) Steady shear viscosity
(c) Viscosity evolution for $\alpha=1$
Figure 3: Thixotropic model validation for a simple shear flow for $\alpha$ range of (1,2,4)). Model parameters are based on the validation process developed by [62]
4 Case Studies: Thixotropic Multiphase Dynamics from Phase Separation to Microfluidics
Following the full validation of the thixotropic multiphase SDPD methodology, we now apply the model to a series of exploratory case studies. These examples are designed to demonstrate the model’s versatility in capturing complex multiphase dynamics under different physical conditions. Specifically, we examine: (i) proteins liquid–liquid phase separation (LLPS), (ii) the behavior of droplet suspensions, (iii) droplet transport through a periodically constricted channel, and (iv) droplet coalescence within microfluidic devices. These scenarios showcase the model’s capability to handle both fundamental and applied problems involving thixotropic behavior and intricate interfacial interactions.
4.1 Liquid-Liquid Phase Separation (LLPS)
Liquid–liquid phase separation is a wide-spread phenomena occurring in biological processes [44, 46, 91], such as the formation of protein droplets in cells to form membraneless organelles [92, 93], intrinsically disordered proteins phase separation in neurodegenerative diseases [88, 89, 90], to name a few. In LLPS the systems can evolve dynamically from homogeneously mixed protein solutions to heterogeneous multiphase-phase systems. The process is characterized by the formation of solute-rich droplets dispersed in a solute-poor continuous phase, where the microstructural properties of the dispersed phase can change over time due to different physical process (coaservation, gelation, cristalization, etc), leading to a thixotropic behavior. In this context, the thixotropic multiphase model can be used to describe the dynamics of proteins LLPS, where phase separation time scales $\lambda_{\sigma}$ and thixotropic time scales $\lambda_{0}$ are interwined.
We investigate systems containing two type of particles, describing a disperse or ”protein phase” ($p$) and a continuous or ”solvent phase” ($s$), with volume fractions $\Phi_{p}$ and $\Phi_{s}=\Phi_{p}$, respectively. As seen in Figure 4, the system evolves from an initial homogeneous condition to a final steady state, where the protein phase separate into droplets. Conceptually, we can think of the protein phase as a collection of protein aggregates dispersed in a continuous solvent phase. The protein phase is characterized by a critical concentration of proteins that triggers the phase separation process, leading to the formation of protein-rich droplets. The solvent phase contains the majority of the solvent and is responsible for the transport and interaction of the protein droplets. Note that we do not model explicitly the protein molecules but rather consider the protein phase as a thixotropic fluid which viscosity changes as protein aggregates form.
For a given $\Phi_{d}$, we can expected that a characteristic time scale $\lambda_{\sigma}$ for the phase separation process lies between a coarsening time scale $R^{2}/D_{p}$ and the coalescence time scale $R\eta_{p}/\sigma_{o}$ ($R^{2}/D_{p}<\lambda_{\sigma}<R/Ca_{v}$), where $R$ is the mean radius of the droplets, $D_{p}\propto k_{B}T/(\eta h^{2})$ is the diffusion coefficient of the protein phase, and $Ca_{v}=\sigma_{o}/\eta_{p}$ is the capillary velocity[46] with $\eta_{p}$ being the viscosity of the protein phase, and $\sigma_{o}$ is the surface tension between the protein droplets and the solvent phase. The coarsening time scale $R^{2}/D_{p}$ represents the time it takes for droplets to grow by diffusion, while the coalescence time scale $R/Ca_{v}$ represents the time it takes for droplets to merge due to surface tension effects. The thixotropic time scale $\lambda_{0}$, governs the rate at which the protein fluid evolves into highly viscous aggregates. The viscosity of the protein phase is assumed to increase with time, with an initial microstructural scalar parameter $f_{0}=0$, as the protein aggregates form and grow. Thus, $\eta_{\infty}$ is the viscosity of the diluted protein phase in a liquid-like state and $\alpha$ determines the maximum viscosity of the protein phase, when $f=1$. To asses the interplay between the phase separation and thixotropic time scales, we investigate the effect of four different parameters on the LLPS process, named (i) protein phase volume fraction $\Phi_{d}$, (ii) capillary velocity $Ca_{v}=\sigma_{o}/\eta_{s}$, (iii) constitutive constant ${1}/{1+\alpha}$, and (iv) thermal energy $k_{B}T$.
Figure 4: Phase separation evolution from the initial homogeneous state to the final phase-separated state.
We conduct LLPS simulations in square domains of size $L=10$, in quiescent conditions ($\dot{\gamma}=0$) such that not microstructural breaking is considered. We use a viscosity of the solvent phase $\eta_{s}=1$ and thixotropic transient values for the protein phase. For the protein volume fraction we consider a range $\Phi_{d}=[5\%,12.5\%,25\%,35\%]$, whereas for the capillary velocity $Ca_{v}$, we explore the range $\{0.1,0.5,1,2\}$. To study the effect of the ratio initial-to-final viscosity $\frac{1}{1+\alpha}$ of the protein phase, we use the range $\frac{1}{1+\alpha}=[0.01,0.05,0.1,0.5]$. Finally, to account for the thermal fluctuations we vary the thermal energy in the range $k_{B}T=[0,0.05,0.1]$. To quantify the impact of these variations, we examine both the number $N_{\text{drop}}$ and size $M_{\text{drop}}$ of the droplets formed upon system stabilization, where ”size” refers to the number of particles comprising each droplet. For each type of test, we perform five independent simulations with different random seeds to ensure statistical robustness. To streamline the analysis and comparison between systems, we compute the Probability Density Function (PDF) of the droplets size, and introduce a set of relevant metrics, such as the average molecular weight (AM), the median (MED), the weighted-average molecular weight (WA), and polydispersity index (PDI). These metrics are defined as follows
$AM=\frac{\sum_{i}^{N_{\text{drop}}}N_{i}M_{i}}{\sum_{i}N_{i}},\quad WA=\frac{\sum_{i}^{N_{\text{drop}}}N_{i}M_{i}^{2}}{\sum_{i}N_{i}M_{i}},\quad PDI=\frac{Mw}{Mn},$
where $N_{\text{drop}}=\sum N_{i}$ and $M_{i}$ is the size of the $i$-th droplet. A value of $PDI=1$ is an indicative of balanced distribution and homogeneity. In turn, values of $PDI>1$ are linked with high dispersion and greater heterogeneity.
In Figure 5(a), we compile the different Kernel Density Estimation (KDE) curve obtained for the different volume fractions. An snapshot of the final state for two concentration is included, where droplets have different colors for clarity. In B (Figure 13(a) and Figure 13(b)), we provide an example of droplet sizes histograms for $\Phi_{d}=5\%$ and $\Phi_{d}=35\%$, illustrating the central tendency statistics.The results indicate that at the lowest volume fractions ($\Phi_{d}=5\%$) investigated, interfacial tension effects are not sufficient to drive the formation of well-defined droplets. In contrast, at higher volume fractions ($\Phi_{d}=35\%$), the protein phase exhibits a tendency to form regular circular droplets, driven by interfacial tension phenomena. Overall, at lower proteins content lead to greater heterogeneity (broader droplet size), whereas as the volume increases, the droplet size distribution tend exhibit lower dispersions.
(a) Protein phase concentration
(b) Capillary velocity
(c) Constitutive constant $\frac{1}{1+\alpha}$
(d) Thermal fluctuations $k_{B}T$
Figure 5: KDE comparison for the different properties analyses including the PID value.
The capillary velocity ($Ca_{v}$) is a property that, as described by Elbaum-Garfinkle et al. [47] and Mitrea et al. [48], is derived from the relationship between surface tension $\sigma_{0}$ and solvent viscosity $\eta_{s}$, mathematically stated as $Ca_{v}=\sigma_{0}/\eta_{s}$. It exhibits linear behavior in relation to the relaxation time and to the characteristic length of the interacting phases. In Figure 5(b), we apply the same statistical analysis for different values of the capillary velocity, namely $Ca_{v}=[0.1,0.2,1,2]$. To facilitate interpretation, histograms of the cluster size distribution and central tendency statistics are provided for $Ca_{v}=0.1$ and $Ca_{v}=2$ in B (Figure 13(c) and Figure 13(d), respectively). These results reveal that for low $Ca_{v}$ values, clusters exhibit irregular shapes and poorly defined structures, resulting in a broad and heterogeneous size distribution skewed toward smaller clusters. In contrast, higher $Ca_{v}$ values promote the formation of well-defined and uniformly sized clusters, driven by the balance between viscosity and surface tension. Overall, low $Ca_{v}$ values are associated with increased heterogeneity in cluster size, whereas higher $Ca_{v}$ values result in more homogeneous and narrowly distributed cluster populations.
In Figure 5(c), we calculate the KDE curves obtained for different values of the constitutive parameter, expressed as $1/(1+\alpha)$, with values $[0.01,0.1,0.2,0.5]$. Representative histograms and central tendency statistics for $0.011/(1+\alpha)=$ 0.01 and $0.51/(1+\alpha)=$ 0.5 are shown in B (Figure 13(e) and Figure 13(f), respectively). The results indicate that, for higher values of $1/(1+\alpha)$, the clusters exhibit well-defined structures with a more uniform and narrowly distributed size profile. Conversely, lower values of $1/(1+\alpha)$ result in poorly developed structures and broader, less concentrated size distributions. Overall, the data show that as $1/(1+\alpha)$ increases, the system becomes more homogeneous, whereas lower values lead to greater heterogeneity in cluster size.
In Figure 5(d), we compile the KDE curves obtained for different values of the thermal fluctuation parameter, $k_{B}T=[0,0.05,0.1]$, as introduced through the stochastic term in the SDPD equations (2). Representative histograms and central tendency statistics for $k_{B}T=0.05$ and $k_{B}T=0.1$ are shown in Appendix B (13(g) and Figure 13(h), respectively). To allow comparison, we run all simulations for 25 time units, as the system would otherwise converge toward a single, large cluster encompassing the entire dispersed phase. In the absence of thermal fluctuations ($k_{B}T=0$), and under conditions where $Ca$ and $\alpha$ do not significantly affect cluster formation, the system exhibits fully developed, circular structures with relatively small and uniform cluster sizes. As thermal fluctuations increase, clusters become less defined and more variable in size, depending on the time evolution. Higher values of $k_{B}T$ accelerate the merging process, eventually leading to a single dominant cluster. This behavior is reflected in the KDE curves, which show that decreased thermal fluctuations result in higher polydispersity and greater heterogeneity in cluster size before coalescence occurs.
4.2 Thixotropic flows in channels: Emulsions and suspensions
The behavior of rigid particles suspended in a fluid was studied by Einstein [51], where the expression $\eta_{eff}=\eta(1+B\phi)$ was derived. This indicates that for concentrations less than 1 ($\phi<<1$) the effective viscosity of the particle suspension is directly proportional to the viscosity of the fluid and the volume fraction of rigid particles, with a constant value $B=2.5$. For deformable particles, Taylor [54] proposed:
$\eta_{eff}=\eta\left[1+\left(\frac{D+2}{2}\right)\phi\left(\frac{\eta^{\prime}+0.4\eta}{\eta^{\prime}+\eta}\right)\right]$ (18)
where $\eta^{\prime}$ is the viscosity of the dispersed phase and $D$ is the system dimension. When $\eta^{\prime}=\eta$, this simplifies to $\eta_{eff}=\eta(1+1.75\phi)$. This model assumes moderate droplet deformability, i.e., low Capillary numbers. We use this expression to validate our multiphase model for emulsions with droplets. We conduct simulations in a square channel of size $L=10$ under simple shear flow $\dot{\gamma}=0.0166$, with particle density $n_{eq}=25$ and no thermal fluctuations. The continuous phase is Newtonian with $\rho=1$, $\eta=15$ and inlet surface tension $\sigma_{0}=2$. Droplets have radius $r_{0}=0.8$ and viscosities $\eta^{\prime}=[15,30]$ covering two viscosity ratios of $\lambda=[1.2]$. This leads to $Re=1.5\times 10^{-3}$ and $Ca=0.1$. We perform a total of eleven simulation with volume fractions in the range $0.30<\phi<$ 0.3 by adjusting the number of droplets. The effective viscosity is calculated by direct measurement of the force on the wall. Figure Figure 6(a) shows the results, including snapshots for $\phi=[0.02,0.14,0.29]$, illustrating droplets deformation. The model captures the increasing trend in viscosity, followed by a slight drop for volumen fraction exceeding 0.2, consistent with prior observations [55, 56, 57, 58].
Once we validate the method for Newtonian cases, we analyze thixotropic flows. Here, the continuous phase has a variable viscosity governed by a thixotropic model, while the dispersed phase retains a constant Newtonian viscosity. We use a channel of length $L=10$, set the continuous phase viscosity to $\eta=1$, and choose thixotropic parameters $\alpha=4$, $\lambda_{0}=100$, and $\beta=1$. These satisfy the condition $\lambda_{0}>L^{2}\rho/\eta_{\infty}$ to guarantee a scale separation between the long microstructural time scale and the short viscous ones [62]. We simulate a simple shear flow ($\dot{\gamma}=0.5$) where a thixotropic continuous phase carries circular dispersed droplets, similar to the Newtonian case. The dispersed phase has a viscosity of $\eta^{\prime}=5$. We test five volume fractions and monitor the emulsion behavior over time. First, we run the system in its Newtonian state until it reaches a steady condition, then switch to the thixotropic model. Results are shown in Figure 6(b). We observe a drop from a peak effective viscosity to a lower steady value as thixotropy develops. Higher concentrations lead to higher effective viscosities.
(a) Droplet suspensions
(b) Thixotropic continuous phase
Figure 6: Emulsion experiments. (a) Validation using Taylor‘s expression for two viscosity ratios $\lambda$ and (b) transient viscosity evolution in a channel with Thixotropic continuous phase for different values of $\phi$
Next, we study the behaviour of emulsions subjected to a Poiseuille flow, starting with the single-phase case. Using a channel of height $L=18$, we apply a body force $g_{x}$ that produces a parabolic velocity profile with a peak value of 6 (in system units). Boundary conditions follow those used in the dynamic validation section. We consider two cases. In the first, we simulate a one-phase flow that begins as Newtonian ($\eta=1$). After reaching steady state, we switch to a thixotropic type with parameters $\alpha=4$, $\lambda_{0}=100$, and $\beta=1$. For comparison, we also run a separate simulation using the maximum viscosity to obtain the limiting velocity profile. Figure 7(a) shows the velocity at $t=100$ (Newtonian state, solid black line) and $t=150$ (thixotropic state, minimum profile). As expected, the velocity decreasesmodel over time as the viscosity increases toward the channel center, reflecting the inverse relationship between viscosity and shear rate in thixotropic fluids.
The second case analyzed has the same objective but is performed for a Poiseuille flow with two phases A and B. We apply the same treatment to a simulation that starts in a Newtonian state and then undergoes transition to the thixotropic model. In this case, phase A maintains its Newtonian behavior throughout the entire simulation, while phase B becomes thixotropic with parameters $\alpha=10$, $\lambda_{0}=100$, and $\beta=1$. As in the previous case, we perform a separate simulation to observe the velocity profile obtained with the maximum viscosity. Figure 7(b) shows the transition between the maximum and minimum velocity profiles. We observe that, as time progresses, the characteristic profile for multiphase flows gradually emerges, as described earlier in the validation section.
(a) One phase
(b) Two phases
Figure 7: Velocity profiles in (a) one and (b) two phases flow. Transition from steady state (both Newtonian fluids) to thixotropic behavior. For the multiphase case only the bottom phase is thixotropic, the upper phase remains Newtonian.
Once we analyze the simple thixotropic cases, we proceed to study emulsions containing droplet suspensions undergoing a Poiseuille flow. For these cases, we use a box of length and height $L=20$ with an array of 25 equidistant and initially undeformed droplets. We consider the continuous phase as Newtonian with viscosity $\eta=5$ and density $\rho=1$, while the dispersed phase exhibits thixotropic behavior with parameters $\eta_{\infty}=5$, $\alpha=4$, $\beta=100$, and $\lambda=100$. We apply the respective body force on the continuous phase to generate the corresponding Poiseuille flow. We analyze two particular cases: one with the formed microstructure, i.e., $f_{0}=1$, and one with the unformed microstructure, $f_{0}=0$. The results appear in the top and bottom rows of Figure 8, respectively. We observe that droplet deformation is smaller in the center of the channel and increases closer to the walls due to the magnitude of the shear rate. Additionally, droplet deformation is smaller in the case with formed microstructure ($f_{0}\approx 1$) since it starts immediately with the highest viscosity and consequently offers a larger hydrodynamic resistance to flow. Conversely, in the case with unformed microstructure ($f_{0}<<1$), the deformation is larger because it starts with the lowest viscosity and evolves over time until reaching its maximum value.
Figure 8: Emulsion with Newtonian continuous phase and thixotropic dispersed phase. Comparison starting with the formed microstructure (top row) versus the unformed microstructure (bottom row).
4.3 Droplet dynamics in a periodically constricted channel
To further investigate the behavior of thixotropic fluids, we analyze the transition of a droplet through a periodically constricted channel. The experiment focuses on observing the changes in viscosity during the contraction and subsequent relaxation of the droplet as it moves through the chamber, driven by the abrupt variations in shear rate between different regions of the domain. To correctly analyze the thixotropic phenomenon it is necessary to define the time scales. As indicated by Vázquez-Quesada et al [60], the sonic time $\tau_{s}=R/c$ which is referred to the time in which a sound wave propagates over a distance $R$ at a sound velocity $c$, must be less than the viscous time ($\tau_{visc}$) to satisfy incompressibility. To determine the viscous relaxation time of the system, we analyze the time decay of the characteristic velocity of the fluid. Considering that in processes dominated by viscous dissipation the velocity follows a decreasing exponential behavior of the form $v(t)\propto\exp(-t/\tau_{visc})$, we calculate $\ln(v)$ as a function of time, thus obtaining a linear relationship. From the linear fit of these data, we relate the slope $m$ of the line to the relaxation time as $\tau_{visc}=-1/m$. These two time scales must meet the condition:
[…]
where $T=L/U$ is the fluid characteristic time. The characteristic thixotropic time scale $\lambda_{0}$ is chosen here on the same order of the characteristic flow time $T$, i.e., $\tau_{s}\ <<\ \tau_{visc}\ <<\ \lambda_{0},\ T$.
We employ for the simulations a channel with periodical boundary conditions of total length 32. Each of the two cross sections has length $L=16$ and a height of the constrained section $H_{c}=2.2$ and for the relaxed section $H_{r}=10$. We consider the transporting fluid to be Newtonian with viscosity $\eta=1$. We treat the droplet as thixotropic with properties $\eta_{\infty}=10$ and $\alpha=4$. We vary the thixotropic parameters $a$ and $b$ to explore scenarios that involve changes in both the characteristic time $\lambda_{0}$ and the structure destruction parameter $\beta$. The density of both the fluid and the droplet is $\rho=1$. The maximum velocity reached by the fluid in the relaxed zone is $U=0.017$ in system units, and the associated sound velocity is $c=20>>U$. We run the simulation long enough to allow the droplet to cycle through the periodic channel several times and stabilize the effective viscosity. With the values described above, we obtain characteristic times of $\tau_{s}=0.8$, $\tau_{visc}=10$, and $T=200$, which comply with the relationship described in Equation 19. We first study the effect of varying the thixotropic time scale by covering four characteristic values $\lambda_{0}=[100,200,400]$, which are on the same time scale as the characteristic time of the fluid $T$. We analyze the viscosity evolution starting with both unformed ($f_{0}=0$) and formed ($f_{0}=1$) microstructure configurations. These results appear in Figure 9(a) and 9(b), respectively. In all the graphs, we normalize the viscosity using the limiting viscosity $\eta_{\infty}$ of the droplet, and we normalize the time axis using the expression $t^{\prime}=t\ \eta_{\infty}/(\rho L^{2})$. Figure 9(a) also shows a detail of the droplet when it is in the constricted zone and its respective relaxation in the channel zone with the largest area.
(a) Initially unformed microstructure, ($f_{0}=0$)
(b) Initially fully formed microstructure ($f_{0}=1$)
Figure 9: Transient evolution of the viscosity of a thixotropic droplet passing through a chamber with diameter change. Comparison at different values of $\lambda_{0}$ at $\beta=1$ for cases starting with a) initially unformed microstructure $f_{0}=0$ and b) initially fully formed microstructure $f_{0}=1$.
For the case of initially unformed microstructure ($f_{0}=0$), it can be seen how the viscosity starts at its limiting value $\eta_{\infty}=10$ and evolves toward its maximum value defined by the thixotropic parameter $\alpha$. This evolution occurs more rapidly when the time scale is smaller, requiring fewer droplet cycles within the channel to reach a periodic stabilization of the viscosity. As $\lambda_{0}$ increases, more cycles are required, and the droplet may even fail to stabilize within the simulated cycles for the highest value of 400. For the case with initially formed microstructure ($f_{0}=1$), it can be seen how the viscosity starts at its maximum peak and decreases until it reaches a periodic stabilization. In this case, all studied values of $\lambda_{0}$ successfully achieve stabilization within the simulated cycles. It is clear that the viscosity decreases in the constriction zone and increases in the relaxed zone due to the abrupt change in the shear rate between these regions. We also investigate the effect of microstructure creation and destruction by varying the parameter $\beta$. We explore four values of $\beta=[0.1,1,10,100]$ for both unformed and formed microstructure. For these simulations, we set the characteristic time scale to $\lambda_{0}=200$, as this value aligns with the characteristic time of the fluid $T$. The results are shown in Figure 10(a) and 10(b), respectively. It can be seen that the direct effect of $\beta$ is to amplify the variations in shear rate and, consequently, the variations in the effective viscosity of the droplet. As the value of $\beta$ decreases, the influence of the shear rate on viscosity diminishes, leading to negligible viscosity variations for $\beta=0.1$ and to maximum variations for $\beta=100$.
(a) Initially unformed microstructure, ($f_{0}=0$)
(b) Initially fully formed microstructure ($f_{0}=1$)
Figure 10: Transient evolution of the viscosity of a thixotropic droplet passing through a chamber with diameter change. Comparison at different values of $\beta$ at $\lambda_{0}=200$ for cases with a) initially unformed microstructure $f_{0}=0$ and b) initially fully formed microstructure $f_{0}=1$.
4.4 Droplet merging using micro-devices
Within microfluidic analysis, the process of droplet merging is of great interest. Various systems have been developed to measure the conditions and parameters involved in this phenomenon. Basu [50] developed a strategy known as Droplet Morphometry and Velocimetry (DMV), a video processing software that enables real-time observation of droplet behavior on small scales, particularly during merging events. Niu et al. [49] extended this research by employing microfluidic devices. To promote merging, they used a system of passive structures (pillars) inside a microfluidic channel, separated by specific distances, which caused droplet deceleration and eventual contact-induced coalescence. Using the SDPD multiphase methodology developed in our study, we numerically simulate this type of complex droplet interaction. Figure 11 shows the initial configuration for the merging process. The physical boundaries (in gray) represent the channel walls and pillar structures. The continuous phase (in white) flows through the channel, while two droplets are initially located at the entrance and exit of the system. We impose periodic boundary conditions in the $x$-direction so that as the final droplet reaches the left boundary of the domain, it reenters from the right as a new incoming droplet. We generate flow within the system by applying a body force. We analyze three particular cases. In the first case, both the penetrating droplet (orange in Figure 11) and the recipient droplet (purple) are Newtonian fluids with equal viscosity $\eta=0.1$. In the second case, the penetrating droplet is thixotropic with properties $\eta_{\infty}=0.1$, $\alpha=9$, $\lambda_{0}=100$, and $\beta=1$, while the recipient droplet remains Newtonian. In the third case, we reverse this configuration, considering the penetrating droplet as Newtonian and the recipient droplet as thixotropic with the same thixotropic parameters. In all cases, we consider the continuous phase as a Newtonian fluid with viscosity $\eta=0.1$. Figure 11 illustrates the merging dynamics. The first droplet enters the chamber and experiences significant resistance due to its interaction with the pillar geometry. Meanwhile, the second droplet advances and begins to penetrate the first droplet, which is partially trapped within the pillar structure. During this interaction, the second droplet transfers part of its kinetic energy to the first, facilitating the joint passage of both droplets through the pillar region. Upon exiting the interaction zone, it can be seen that the smaller droplet is almost fully embedded within the larger one. This interaction results from surface tension, which promotes coalescence and stabilizes the new droplet structure formed after fusion. Figure 11 also shows the final merged droplet for each studied case. It can be seen that in Case 1 (Newtonian/Newtonian), the penetrating droplet is nearly fully incorporated into the recipient droplet, resulting in a poorly defined internal structure with dispersed particles. In Case 2 (thixotropic/Newtonian), we observe a similar pattern, though with slightly more organized particles. Case 3 (Newtonian/thixotropic) exhibits a notably different result, with the formation of a flat and well-defined interface between the two droplets. This distinct boundary reflects the strong resistance imposed by the thixotropic recipient phase, which significantly slows down the penetration of the Newtonian droplet. As a consequence, the internal particle arrangement is more structured, clearly indicating the influence of the rheological properties of the recipient droplet on the merging dynamics.
Figure 11: Initial state and merging process using microfluidic device. For Case 1, both droplets were considered to be Newtonian. In Case 2 the penetrating droplet (orange) is thixotropic and the recipient droplet (purple) is Newtonian. In Case 3 the penetrating droplet is Newtonian and the recipient droplet is thixotropic.
5 Conclusions
We proposed a numerical methodology based on Smoothed Dissipative Particle Dynamics (SDPD) for simulating multiphase flows with thixotropic behavior. The model accounts for surface tension and tracks the time-dependent evolution of viscosity due to microstructural changes. The approach was validated through static and dynamic cases, including droplet deformation, contact angles, Poiseuille flow, and flow past obstacles. It accurately reproduced known behaviors and matched analytical predictions, confirming its reliability.
The validated thixotropic multiphase SDPD model was applied to a range of complex flow scenarios, demonstrating its potential for investigating biologically and industrially relevant systems. In the case of liquid–liquid phase separation (LLPS), the model captured how parameters such as protein concentration, capillary number, and microstructural relaxation influence droplet morphology—revealing conditions that favor the formation of stable, well-defined aggregates. Simulations of suspensions and emulsions showed that the model accurately reproduces the rheological evolution of complex fluids under confinement and shear, including effective viscosity trends and interfacial dynamics in two-phase systems. In microfluidic geometries, such as periodically constricted channels and droplet merging devices, the model captured key behaviors like localized viscosity variation, flow stabilization, and coalescence resistance due to internal structure. These results highlight the model’s capacity to simulate time-dependent rheology and interface behavior, making it a promising tool for the design and analysis of systems in soft matter physics, biomedical applications, and microfluidic device engineering.
Appendix A Simple Newtonian Flows
We carried out the dynamic validation of the model considering several scenarios. The first part of the validation involves the development of a Poiseuille flow for a single-phase flow ($\alpha$) in three variations: simple profile, reverse Poiseuille and flow around a cylinder. For particle methods, as indicated by Backer et al. [61], the average velocity and shear stress in a Poiseuille flow are given by
[…]
$\tau_{xy}$ $=\rho g_{x}\left(y-\frac{1}{2}L\right),$ (21)
where $\rho$ is the density, $L$ the length between the parallel walls and $g_{x}$ the body force. For the simple and reverse case, we employ a two-dimensional square channel with walls separated by $L=10$, $\rho=1$ and $\eta=5$ in the units system. We select the equilibrium particle density $n_{eq}=25$ and we do not taken into account the effects of thermal fluctuations. We compare three velocities under laminar flow regime determine by the Reynolds number $Re=\rho v_{x}D/\eta$. We configure the body force with the required values to obtain the respective predefined velocities. In the case of reverse Poiseuille flow, the body force is positive in one half of the channel and negative in the opposite half, hence resulting in the associated representative velocity profile. We choose the speed of sound to be at least 20 times greater than the maximum velocity reached by the flow. We established a periodicity condition along the x-direction and we define a no-slip condition for walls by setting the related particles as totally stationary and without any contribution to the fluid. Figures 12(a), 12(b), 12(d) and 12(e) shows the velocity and shear profiles calculate by our model in comparison to the theoretical values obtain by applying Equation 20 for $Re=[0.1,0.5,1]$. Regarding the case of flow around a cylinder, we perform a validation based on the numerical results obtained by Ellero and Adams [40] using both SPH and immersed boundary method (IBM). We employ a channel characterized by a centrally located cylinder with radius $R_{c}=2$ and height $L=4R_{c}$ to study the flow characteristics. Considering the condition of periodicity on the x-direction, we set the width of the channel $L_{c}=6R_{c}$ such that the distance between repeated cylinders allows the fluid development. For this validation, the same physical parameters used in the first two cases are kept with the only exception of the viscosity whose new value is $\eta=1$ (in line with the approach of Ellero and Adams [40]). We configure the cylinder as well as the channel walls with a no-slip condition. Figures 12(c) and 12(f) shows the velocity profiles versus the normalized $x/R_{c}$ and $y/R_{c}$ axis for both for the results obtained with our approach and those getting by the IBM [40]. These curves are taken on a vertical plane at a distance of $x/R_{c}=3$ and a horizontal plane at a distance of $y/R_{c}=3.5$, respectively. As can be noted for the three analyses cases of single-phase flow, the comparative results between the theoretical values and the numerical approximations exhibit a close agreement showing the accuracy of the model in order to capture the characteristics of the flow under dynamic conditions.
(a) Simple Poiseuille velocity profiles
(b) Reverse Poiseuille velocity profiles
(c) Flow around a cylinder. Section A.
(d) Simple Poiseuille shear stress profiles
(e) Reverse Poiseuille shear stress profiles
(f) Flow around a cylinder. Section B
Figure 12: Validation of the methodology for one phase Poiseuille flows using (a)-(d) simple profile, (b)-(e) reverse profile and (c)-(f) flow around a cylinder in a vertical and horizontal plane.
Appendix B Droplet size distribution
We performe numerical simulations using the multiphase SDPD methodology proposed in this work for the analysis of LLPS phenomena and examined four main properties: (i) protein phase volume fraction $\Phi_{d}$, (ii) capillary velocity $Ca_{v}=\sigma_{o}/\eta_{s}$, (iii) constitutive constant ${1}/{1+\alpha}$, and (iv) thermal energy $k_{B}T$. For each property we construct histograms and estimate the probability density using the KDE function. In addition, we calculate the arithmetic mean, the median and the Weight Average Molecular Weight. In this appendix we present the histograms corresponding to each property, accompanied by a representative snapshot of the last simulated time. The results for protein phase volume fraction $\Phi_{d}=[5\%,35\%]$, $Ca_{v}=[0.1,2]$, ${1}/{1+\alpha}=[0.5,0.01]$ and $k_{B}T=[0.05,0.1]$ are shown in Figure 13.
(a) Protein phase concentration $\Phi_{d}=5\%$
(b) Protein phase concentration $\Phi_{d}=35\%$
(c) Capillary velocity $Ca=0.1$
(d) Capillary velocity $Ca=2$
(e) Constitutive constant $\frac{1}{1+\alpha}=0.5$
(f) Constitutive constant $\frac{1}{1+\alpha}=0.01$
(g) Thermal fluctuations $k_{B}T=0.05$
(h) Thermal fluctuations $k_{B}T=0.1$
Figure 13: Cluster histograms for the comparison of (a) protein phase concentration, (b) capillary velocity, (c) constitutive constant $\alpha$ and (d) thermal fluctuations.